# Characterizing Pattern Matching and Its Limits on Compositional Task Structures

### v2

@@ F.3 Backgrounds, Useful Facts, and Lemmas
We also prove the tail probability bound for the coupon collector’s problem (newman1960double; erdHos1961classical) here. We again restate the lemma here for the sake of readability.
@@
Poissonization (kac1949deviations; holst1986birthday; aldous1989probability; jacquet1998analytical; johansson1998longest; borodin2000asymptotics) can be thought of as a special case of randomization using Poisson random variables. A Poisson random variable $N\sim\mathrm{Poisson}(\lambda)$ with a parameter $\lambda>0$ is equipped with a probability mass function
@@
The only things left to obtain a complete (de)-Poissonization lemma for upper/lower-bounding the sequence $a_{n}$ are the tail probability bounds for a Poisson distribution. The concentration inequalities for the Poisson distribution are already well-known, [^8] For example, see a document online: “A short note on Poisson tail bounds.” although we derive slightly different forms of them for our own purposes.
@@
Moreover, it is sufficient for proving the first equation to study the case with non-negative integers $z_{1},\cdots,z_{m}$ such that $\sum_{i=1}^{m}z_{i}=n$; otherwise, both sides of Eq.˜17 are zero. In this case, observe an inclusion between events
@@
Therefore, we have just proved Eq.˜17.
@@
This proves Eq.˜19, concluding the proof. ∎
@@ F.3.2 Binomial Random Intersection Graphs
Let us explain the random $k$-intersection graph (singer1995random; karonski1999random; godehardt2003two). Consider a set $V$ of vertices and another set $W$ of items. Each vertex $v\in V$ is randomly assigned a subset of items $W_{v}\subset W$. Then, a pair of vertices $u$ and $v$ are adjacent (i.e., connected with an edge) in a random $k$-intersection graph if and only if at least $k$ items are shared between $u$ and $v$, i.e., $\left|W_{u}\cap W_{v}\right|\geq k$. We remark that $k$ is a universal constant throughout our paper.
@@
In addition, observe that the set $S(x_{12},x^{\prime}_{12}\mid D)$ defined in Eq.˜6 is identical to the intersection $W_{x_{12}}\cap W_{x^{\prime}_{12}}$. Hence, by the definition of the edge set $E_{b}^{(D,k)}$ in Eq.˜5, two distinct vertices $x_{12}$ and $x^{\prime}_{12}$ are adjacent if and only if $\left|W_{x_{12}}\cap W_{x^{\prime}_{12}}\right|\geq k$. In summary, every $b$-evidence graph ${\mathcal{G}}_{b}^{(D,k)}$ is a binomial random $k$-intersection graph ${\mathcal{G}}^{(k)}(n_{b},m,p)$ with parameters $n_{b}=\left|V_{b}\right|$, $m=\left|{\mathcal{X}}_{3}\right|$, and $p=1-\exp(-\lambda/\left|{\mathbb{X}}\right|)$. ∎
@@
Here, we review a few of the seminal results on the connectivity (as well as the disappearance of isolated vertices) of binomial random $k$-intersection graphs (singer1995random; zhao2014k; zhao2017connectivity; rybarczyk2011sharp; rybarczyk2017coupling). [^9] Caveat: in the literature of (random) graph theory, the letter ‘$k$’ is often used for $k$-connectivity (i.e., being connected after removing less than $k$ vertices/edges). On the other hand, they study ‘random $s$-intersection graphs,’ using $s$ as the intersection constraint parameter. The proofs are involved; thus, we omit them.
@@ G.1 Measurement Protocol for $N_{\mathrm{req}}$
Specifically, our criterion for “reliable generalization” on ID is defined as:
  • The model must reach ID test accuracy of 0.99 within 100 epochs after achieving training accuracy > 0.99.
This protocol balances several considerations:
@@
    Measurement precision: For each identified $N_{\mathrm{req}}$, we verified that 75% of this dataset size consistently failed to meet our generalization criterion. This establishes that our measurement error is at most $\log(0.75)=0.125$ in log scale, providing confidence in the derived power-law exponents.
@@ G.2 Measured Power-Law Scaling Exponents Across Task Structures and Model Sizes
Table 2 presents the power-law exponents obtained by linear fitting $\log(\left|{\mathcal{X}}\right|)$ vs. $\log(N_{\mathrm{req}})$ plots, all with $R^{2}>0.99$. The consistency of exponents across model sizes suggests that the observed power-law scaling relates to properties of the compositional tasks themselves, rather than model capacity. This observation aligns with our theoretical derivation in Section 5.1, which predicts that the required dataset size scales at least quadratically with token set size.
@@ G.3 Robustness to Hyperparameter Variations
To verify that our observed power-law scaling relationship is not an artifact of specific hyperparameter choices, we conduct ablation studies with modified training configurations. Figure 16 demonstrates that for the 2-Hop task with $\left|{\mathcal{X}}\right|=50$, the following changes did not significantly affect the measured $N_{\mathrm{req}}$ or the derived power-law exponent:
@@
    Learning rate reduction: Halving the learning rate from 8e-4 to 4e-4
@@
    Weight decay reduction: Decreasing weight decay by a factor of 10 (from 0.1 to 0.01)
@@
    Generalization criteria modification: Requiring test accuracy > 0.95 within 10 epochs after training accuracy > 0.95
@@ Appendix I Detailed Discussion on the Taxonomy for Understanding Generalization Mechanisms
Type-II: Function property-based generalization. This mechanism exploits intrinsic properties of individual primitive functions, e.g., algebraic invariances such as commutativity or input irrelevance, where certain arguments never affect the output (e.g., $f(x_{1},x_{2})=f(x_{1})$ even when distractor $x_{2}$ is present (wen2024sparse)). Unlike the previous type, this mechanism explains the generalization beyond the coverage by leveraging ‘global’ properties that hold across all possible inputs of a primitive, beyond what is actually observed. We interpret the Reversal Curse phenomenon (berglund2023reversal) as an example of the layered nature of challenges across multiple generalization types. Our framework predicts the failure of pattern matching on this problem, since training on “$A$ is BB” provides no functional equivalence evidence for “$B$ is $A^{-1}$”. An architectural modification to learn inverse mappings from the same training data to handle this problem (lv2024analysis) can be interpreted as a utilization of Type-II generalization to enable generalization beyond coverage.
Type-III: Shared-operator generalization. This mechanism emerges through the reuse of identical primitive functions across computational positions (e.g., when $f_{1}=f_{2}$). Recurrent architectures (hochreiter1997long) exemplify the utilization of this through weight sharing across time steps, enabling processing of variable-length sequences (graves2014neural). Similarly, it has been reported in Transformers with inductive biases towards reuse of the same computation through parameter sharing (dehghani2018universal; csordas2021devil; wang2024grokked) can improve generalization on complex compositional tasks where the same primitive function can be reused in various contexts. We interpret this mechanism as exploiting structural repetition.
@@
Compared to prior categorizations of generalization, which focus on observed phenomena (lake2018generalization; hupkes2020compositionality), we categorize the underlying mechanisms. As noted in Sec.˜1, many behavioral studies have examined tasks mixing functional equivalence, primitives’ intrinsic properties, and operator reuse within the same benchmark, making it difficult to pinpoint the true source of success or failure. We therefore advocate clearer experimental control and community discussion around this mechanistic distinction to sharpen future analyses of neural generalization.

### v3

@@ F.3 Backgrounds, Useful Facts, and Lemmas
We also prove the tail probability bound for the coupon collector’s problem (Newman, 1960; Erdős and Rényi, 1961) here. We again restate the lemma here for the sake of readability.
@@
Poissonization (Kac, 1949; Holst, 1986; Aldous, 1989; Jacquet and Szpankowski, 1998; Johansson, 1998; Borodin et al., 2000) can be thought of as a special case of randomization using Poisson random variables. A Poisson random variable $N\sim\mathrm{Poisson}(\lambda)$ with a parameter $\lambda>0$ is equipped with a probability mass function
@@
The only things left to obtain a complete (de)-Poissonization lemma for upper/lower-bounding the sequence $a_{n}$ are the tail probability bounds for a Poisson distribution. The concentration inequalities for the Poisson distribution are already well-known, [^8] For example, see a document online: “A short note on Poisson tail bounds.” although we derive slightly different forms of them for our own purpose.
@@
Moreover, it is sufficient for proving the first equation to study the case with non-negative integers $z_{1},\cdots,z_{m}$ such that $\sum_{i=1}^{m}z_{i}=n$; otherwise, both sides of Eq.˜16 are zero. In this case, observe an inclusion between events
@@
Therefore, we have just proved Eq.˜16.
@@
This proves Eq.˜18, concluding the proof. ∎
@@ F.3.2 Binomial Random Intersection Graphs
Let us explain the random $k$-intersection graph (B., 1995; Karoński et al., 1999; Godehardt and Jaworski, 2003). Consider a set $V$ of vertices and another set $W$ of items. Each vertex $v\in V$ is randomly assigned a subset of items $W_{v}\subset W$. Then, a pair of vertices $u$ and $v$ are adjacent (i.e., connected with an edge) in a random $k$-intersection graph if and only if at least $k$ items are shared between $u$ and $v$, i.e., $\left|W_{u}\cap W_{v}\right|\geq k$. We remark that $k$ is a universal constant throughout our paper.
@@
In addition, observe that the set $S(x_{12},x^{\prime}_{12}\mid D)$ defined in Eq.˜5 is identical to the intersection $W_{x_{12}}\cap W_{x^{\prime}_{12}}$. Hence, by the definition of the edge set $E_{b}^{(D,k)}$ in Eq.˜4, two distinct vertices $x_{12}$ and $x^{\prime}_{12}$ are adjacent if and only if $\left|W_{x_{12}}\cap W_{x^{\prime}_{12}}\right|\geq k$. In summary, every $b$-evidence graph ${\mathcal{G}}_{b}^{(D,k)}$ is a binomial random $k$-intersection graph ${\mathcal{G}}^{(k)}(n_{b},m,p)$ with parameters $n_{b}=\left|V_{b}\right|$, $m=\left|{\mathcal{X}}_{3}\right|$, and $p=1-\exp(-\lambda/\left|{\mathbb{X}}\right|)$. ∎
@@
Here, we review a few of the seminal results on the connectivity (as well as the disappearance of isolated vertices) of binomial random $k$-intersection graphs (B., 1995; Zhao et al., 2014; 2017; Rybarczyk, 2011; 2017). [^9] Caveat: in the literature of (random) graph theory, the letter ‘$k$’ is often used for $k$-connectivity (i.e., being connected after removing less than $k$ vertices/edges). On the other hand, they study ‘random $s$-intersection graphs,’ using ss—instead of $k$ as this paper—as the intersection constraint parameter. The proofs are involved; thus, we omit them.
@@ G.1 Measurement Protocol for $N_{\mathrm{req}}$
Specifically, our criterion for the “reliable generalization” on ID is defined as: The model must reach ID test accuracy of 0.99 within 100 epochs after achieving training accuracy > 0.99. This protocol balances several considerations:
@@
    Measurement precision: For each identified $N_{\mathrm{req}}$, we verified that 75% of this dataset size consistently failed to meet our generalization criterion. This establishes that our measurement error is at most $\log(0.75)\approx 0.125$ in log scale, providing confidence in the derived power-law exponents.
@@ G.2 Measured Power-Law Scaling Exponents Across Task Structures and Model Sizes
Tab.˜2 presents the power-law exponents obtained by linear fitting $\log(\left|{\mathcal{X}}\right|)$ vs. $\log(N_{\mathrm{req}})$ plots, all with $R^{2}>0.99$. The consistency of exponents across model sizes suggests that the observed power-law scaling relates to properties of the compositional tasks themselves, rather than model capacity. This observation aligns with our theoretical derivation in Section 5.1, which predicts that the required dataset size scales at least quadratically with token set size.
@@ G.3 Robustness to Hyperparameter Variations
To verify that the power-law scaling relationship we observed is not an artifact of specific hyperparameter choices, we conduct ablation studies with modified training configurations. Fig.˜16 demonstrates that for the 2-Hop task with $\left|{\mathcal{X}}\right|=50$, the following changes did not significantly affect the measured $N_{\mathrm{req}}$ or the derived power-law exponent:
@@
    Learning rate reduction: Halving the learning rate from 8e-4 to 4e-4;
@@
    Weight decay reduction: Decreasing weight decay by a factor of 10 (from 0.1 to 0.01);
@@
    Generalization criteria modification: Requiring test accuracy > 0.95 within 10 epochs after training accuracy > 0.95.
@@ Appendix I Detailed Discussion on the Taxonomy for Understanding Generalization Mechanisms
Type-II: Function property-based generalization. This mechanism exploits intrinsic properties of individual primitive functions, e.g., algebraic invariances such as commutativity or input irrelevance, where certain arguments never affect the output (e.g., $f(x_{1},x_{2})=f(x_{1})$ even when distractor $x_{2}$ is present (Wen et al., 2025)). Unlike the previous type, this mechanism explains the generalization beyond the coverage by leveraging ‘global’ properties that hold across all possible inputs of a primitive, beyond what is actually observed. We interpret the Reversal Curse phenomenon (Berglund et al., 2024) as an example of the layered nature of challenges across multiple generalization types. Our framework predicts the failure of pattern matching on this problem, since training on “$A$ is BB” provides no functional equivalence evidence for “$B$ is AA”. An architectural modification to learn inverse mappings from the same training data to handle this problem (Lv et al., 2024) can be interpreted as a utilization of Type-II generalization to enable generalization beyond coverage.
Type-III: Shared-operator generalization. This mechanism emerges through the reuse of identical primitive functions across computational positions (e.g., when $f_{1}=f_{2}$). Recurrent architectures (Hochreiter and Schmidhuber, 1997) exemplify the utilization of this through weight sharing across time steps, enabling processing of variable-length sequences (Graves et al., 2014). Likewise, it has been reported in Transformers with inductive biases towards reuse of the same computation through parameter sharing (Dehghani et al., 2019; Csordás et al., 2021; Wang et al., 2024a) can improve generalization on complex compositional tasks where the same primitive function can be reused in various contexts. With a similar insight, the generalization capability of Transformers in terms of sequence length (i.e., length generalization) and its limitation has recently been studied a lot, especially in terms of arithmetic/algorithmic tasks (Anil et al., 2022; Zhou et al., 2024a; Shen et al., 2023; Zhou et al., 2024b; Cho et al., 2024; 2025). We interpret this mechanism as exploiting structural repetition.
@@
Compared to prior categorizations of generalization, which focus on observed phenomena (Lake and Baroni, 2018; Hupkes et al., 2020), we categorize the underlying mechanisms. As noted in Sec.˜1, many behavioral studies have examined tasks mixing functional equivalence, primitives’ intrinsic properties, and operator reuse within the same benchmark, making it difficult to pinpoint the true source of success or failure. We therefore advocate clearer experimental control and community discussion around this mechanistic distinction to sharpen future analyses of neural generalization.